OpenStudy (anonymous):

Given a wavelength of 121,270 millimeters, what is the wavelength in kilometers

OpenStudy (anonymous):

1km = 10^3m 1m = 10^2cm 1cm = 10^1mm Can you work it out from there?

OpenStudy (anonymous):

Yes thanks!

OpenStudy (anonymous):

No worries =)